OpenStudy (anonymous):

Let f(x) = 3x^2 + 10x - 8 and g(x) = x + 4 FIND ( g of f)(2)

OpenStudy (anonymous):

Lets do this one together.

OpenStudy (anonymous):

Okay, let me know when you want my answer :-)

OpenStudy (anonymous):

Nvm i blanked. ; - ;

OpenStudy (anonymous):

f(2) = 3(2)^2 +10(2) -8 =24 g(24) = (24) +4 =28

OpenStudy (anonymous):

I got six I do it the long way with replacing x and then with the number. ;/

OpenStudy (anonymous):

I'll do it over the way you did it.

sam (.sam.):

gf(2)

OpenStudy (anonymous):

g(3x^2 + 10x - 8) = (3x^2 + 10x - 8) +4 g(f(2)) = (3(2)^2 +10(2) -8) +4 = (12 +20 -8) +4 =28